# A Generalized Back-Door Criterion for Linear Regression

### v3

@@ 1 Introduction
Meanwhile, some previous studies have explored the identification of partial causal relationships. Pearl1998 proposed the single-door criterion for identifying direct effects, and Avin2005 and Shpitser2008 studied the identifiability of path-specific effects. BritoandPearl2002, Tian2009, Foygel2012, Chen2016, and Barber2022 discussed the identifiability of path coefficients by considering the paths that contribute to covariance. In addition to the studies mentioned above, various other previous works exist; however, conditions guaranteeing that a partial regression coefficient coincides with a causal effect not mediated by covariates have been identified only in cases where the direct or total effect is of interest. Consequently, the logical validity of the causal interpretation of partial regression coefficients, as mentioned at the beginning, remains insufficiently established. Given that regression analysis is widely used not only in academic research but also in various practical contexts, it is crucial to provide a theoretically sound basis for its interpretation.
@@
Unlike the direction of generalization taken in previous studies, the selective-door criterion focuses solely on preserving directed paths that remain unblocked by covariates. As a result, it provides a clear framework for identifying partial causal relationships. Moreover, whereas Pearl1998 discussed regression coefficients that minimize the mean squared error under conditional distributions of linear Gaussian models, in this paper, without assuming any specific distribution for the error terms and by employing uncorrelatedness, I present general results for standard partial regression coefficients. This means that linear causal effects can be estimated by the least squares method without assuming independence or normality of the error terms in the data-generating process.
@@ 2.1 Model Setup
where $\bm{PA}(X_{i})\subset\bm{V}\setminus\{X_{i}\}$ denotes the set of immediate causes (parents) of $X_{i}$. Here, $a_{ij}$ is a nonzero constant referred to as the path coefficient from $X_{j}$ to $X_{i}$, and $u_{1},\dots,u_{N}$ are error terms interpreted as aggregations of all causes not included in $\bm{V}$. In studies of linear SEM, error terms are often assumed to be mutually independent or normally distributed. This paper relaxes this assumption, assuming instead that they are uncorrelated.
@@ 2.4 Controlled Total Effects
Using the do-operator Pearl2009, this quantity can be interpreted as
@@
See Example 4.2 in the next section for a concrete illustration of this measure.
@@ 3.2 Main Results
Pearl1998 has shown that these results hold under certain assumptions for conditional regression coefficients, i.e., the linear best approximation coefficients under the conditional distribution. A contribution of the present study is to demonstrate that causal effects can be consistently estimated using ordinary least squares, even in the absence of assumptions regarding the independence or normality of the error terms in the data-generating process.

### v4

@@ 1 Introduction
Meanwhile, some previous studies have explored the identification of partial causal relationships. Pearl1998 proposed the single-door criterion for identifying direct effects, and Avin2005 and Shpitser2008 studied the identifiability of path-specific effects. BritoandPearl2002, Tian2009, Foygel2012, Chen2016, and Barber2022 discussed the identifiability of path coefficients by considering the paths that contribute to covariance. Fulcher2020 studied a generalized front-door criterion for identifying indirect effects. However, conditions guaranteeing that a partial regression coefficient coincides with a causal effect not mediated by covariates have been identified only in cases where the direct or total effect is of interest. Consequently, the logical validity of the causal interpretation of partial regression coefficients, as mentioned at the beginning, remains insufficiently established. Given that regression analysis is widely used not only in academic research but also in various practical contexts, it is crucial to provide a theoretically sound basis for its interpretation.
@@
Unlike the direction of generalization taken in previous studies, the selective-door criterion focuses solely on preserving directed paths that remain unblocked by covariates. As a result, it provides a clear framework for identifying partial causal relationships. Furthermore, I prove the general relationship between partial regression coefficients and graphical criteria by relying on uncorrelatedness rather than on properties specific to linear Gaussian models. Although this approach may render the proof somewhat more verbose, it allows for a clearer demonstration that post-treatment bias can be characterized by a sequence of vertices formed by concatenating a directed path with a back-door path. Since the properties of partial regression coefficients can be proven under specific distributional assumptions without loss of generality, the existing results of Spirtes1998 and Pearl1998 are established by making use of properties of linear Gaussian models. Perhaps for this reason, it does not seem to be widely recognized that the assumption of normality is in fact unnecessary. The proof of the generalized theorem presented in this paper also reaffirms that normality does not need to be assumed.
@@ 2.1 Model Setup
where $\bm{PA}(X_{i})\subset\bm{V}\setminus\{X_{i}\}$ denotes the set of immediate causes (parents) of $X_{i}$. Here, $a_{ij}$ is a nonzero constant referred to as the path coefficient from $X_{j}$ to $X_{i}$, and $u_{1},\dots,u_{N}$ are error terms interpreted as aggregations of all causes not included in $\bm{V}$. We do not assume any specific distribution for the error terms and assume that they are mutually uncorrelated.
@@ 2.4 Controlled Total Effects
Using the do-operator (Pearl2009), this quantity can be “interpreted” as
@@
Since independence among the error terms is not assumed, this equality cannot be derived based on do-calculus(Pearl2009). However, given that the notation involving the do-operator represents probabilistic concepts of the post-intervention distribution, this interpretation is appropriate. See Example 4.2 in the next section for a concrete illustration of this measure.
@@ 3.2 Main Results
These corollaries are originally established in Spirtes1998 and Pearl1998 based on conditional independence relationships and conditional covariances in linear Gaussian models. However, as also noted in Spirtes1998 and Nandy2017, their proofs are without loss of generality, since the partial regression coefficients depend solely on $(\bm{I}_{N}-\bm{A})^{-1}\mathop{\rm Var}\nolimits(\bm{u})(\bm{I}_{N}-\bm{A})^{-1}$. Here, $\bm{I}_{N}$ denotes the $N\times N$ identity matrix. A contribution of the present study is to show that linear causal effects not mediated by covariates can also be consistently estimated using ordinary least squares, even without assuming independence or normality of the error terms in the data-generating process.